Adamson Corporation is considering four average-risk projects with the following costs and rates of return: Project ExpectedAny help would be appreciated - thanks! :)
Answer #1

a. Cost of debt =11%
Cost of Preferred Shares =Dividend/Price =6/56 =10.71429% or 10.71%
Cost of Retained Earnings =D1/Price+growth =4.25/37+4% =15.4865% or 15.49%

b. WACC
=Weight of common stock*Cost of equity+Weight of Preferred Stock*Cost of Preferred Stock+Weight of Debt*Cost of Debt*(1-tax Rate) =75%*15.4865%+10%*10.71429%+15%*11%*(1-30%) =13.84%

c. Project 1 should be accepted as WACC will be higher.
Project 2 should be accepted
Project 3 should be rejected
Project 4 should be rejected.